# Reliable-loc: Robust sequential LiDAR global localization in large-scale street scenes based on verifiable cues

### v1

@@ 6.1 Effectiveness of the spatial verification-based observation model
To further demonstrate the effectiveness of the spatial verification-based observation model, we extracted four 100-frame clips from the experimental data of Jiefang road: 500-600, 1000-1100, 1500-1600, and 2000-2100. We then compared the localization performance of PF-loc, PF-SGV-loc, and PF-SGV2-loc on these clips. The particles are initialized with a known position but an unknown yaw in the experiment. The experimental results are presented in Fig.9.
@@ 6.2 Necessity of switching the localization mode
To illustrate the necessity of switching the localization mode, we select representative clips of lengths 750, 500, 500, and 500 frames on the data Info campus, Zhongshan park, Jiefang road, and Yanjiang road 1 respectively. We use Reg-loc and Reliable-loc to perform global localization in these clips, the position and yaw angle are known at particle initialization, and the localization trajectories are shown in Fig.10.
@@ 6.3 Distribution of the localization mode
width = colspec = Q[140]Q[100]Q[100]Q[110]Q[110]Q[110]Q[110], hlines = 0.08em, Methods & CS college Info campus Zhongshan park Jiefang road Yanjiang road 1 Yanjiang road 2
Reg Mode Ratio (%) 85.61 65.56 90.69 83.5 36.78 87.24
@@ 6.4 Parameter analysis
As shown in Table 4, Reliable-loc exhibits reduced sensitivity to parameters in Info campus, achieving position and yaw accuracies of approximately 2.5 meters and 4 degrees, respectively. For CS college, Zhongshan park, Jiefang road, and Yanjiang road 2, the optimal parameter is 5e-4. For Yanjiang road 1, the optimal parameter is 1e-4. Reliable-loc demonstrates its ability to achieve reliable localization across various parameter settings, maintaining relatively stable performance on the experimental data. Typically, optimal localization performance is achieved when $\lambda_{thr}$ is set to 1e-4, 5e-4, or 1e-3. In practical applications, a setting of 5e-4 for $\lambda_{thr}$ is recommended.
@@
width = colspec = Q[129]Q[119]Q[94]Q[106]Q[131]Q[106]Q[125]Q[125], cell21 = r=2, cell41 = r=2, cell61 = r=2, cell81 = r=2, cell101 = r=2, cell121 = r=2, cell141 = r=2, hline1-2,4,6,8,10,12,14,16 = -0.08em, Param (unit:e-4) & Loc Error CS college Info campus Zhongshan park Jiefang road Yanjiang road 1 Yanjiang road 2
1 Position (m)↓ 0.87 2.56 1.47 1.92 2.39 1.51
Yaw (deg)↓ 2.17 4.19 2.69 2.81 4.45 3.37
5 Position (m)↓ 0.73 2.41 1.13 1.92 3.60 1.33
Yaw (deg)↓ 2.09 4.02 2.44 2.64 5.62 3.22
10 Position (m)↓ 0.90 2.30 4.14 2.18 2.89 2.18
Yaw (deg)↓ 2.53 4.06 3.02 3.24 4.74 3.16
20 Position (m)↓ 0.91 2.38 1.43 2.21 3.06 3.21
Yaw (deg)↓ 2.18 4.20 3.09 3.40 4.21 4.32
40 Position (m)↓ 0.83 2.22 2.27 2.59 3.18 4.01
Yaw (deg)↓ 2.30 3.66 3.73 4.10 4.23 4.77
80 Position (m)↓ 0.78 2.32 2.32 2.75 3.18 3.95
Yaw (deg)↓ 2.49 4.16 3.77 4.06 4.23 4.55
160 Position (m)↓ 1.36 2.33 2.94 2.76 3.18 3.95
Yaw (deg)↓ 2.54 4.07 3.99 4.18 4.23 4.55
@@ 6.5 Efficiency analysis
As shown in Table 5, Reliable-loc requires an average of 68.08ms to process one submap. The MCL-based methods exhibit longer total elapsed times, exceeding Reliable-loc by approximately 100ms. Evidently, Reliable-loc demonstrates superior time efficiency, making it potential for real-time applications.
@@ 6.6 Deficiencies and future work
The above experimental results demonstrate that Reliable-loc can achieve reliable localization in large-scale, complex street scenes by integrating place recognition and MCL, with a position accuracy of up to 0.73m and yaw accuracy of up to 2.09 degrees. Nevertheless, the method’s robustness and accuracy rely on the descriptiveness and discrimination of the global and local features from place recognition. When Reliable-loc starts localization in a scene with extremely insufficient features, the particles are prone to converging to the incorrect region due to the poor discrimination of the global features from place recognition, ultimately resulting in localization failure. A corresponding failure case is shown in Fig.12, and Fig.12 shows the satellite image, helmet point cloud, the curve of the positional error, and the number of grids occupied by the particles over time.
@@
In the future, we will conduct research in three areas to enhance the robustness and accuracy of the localization systems. Firstly, building on existing place recognition efforts, we will introduce contrastive learning to strengthen the descriptiveness, discrimination, and generalization of features extracted by the network, thereby improving the reliability of the localization systems in challenging scenes such as feature insufficiency (Jing and Tian, 2020; Xie et al., 2020; Wu et al., 2023). Secondly, we will explore continuous learning to enable the place recognition model to evolve, overcoming challenges posed by scene switching and sensor changes during long-term localization, and enhancing the system’s resilience (Knights et al., 2022; Yin et al., 2023b). Finally, referencing the backend optimization framework in Simultaneous Localization and Mapping, we will fully leverage the information contained in raw data from multiple sensors like IMU and LiDAR to achieve more reliable and accurate localization in a tightly coupled manner (Chiang et al., 2019; Ye et al., 2019; Pan et al., 2021).
@@ 7 Conclusion
In this paper, we propose a LiDAR-based sequential global localization method that effectively overcomes challenges posed by feature insufficiency and incomplete map coverage by utilizing spatial and temporal verifiable cues, enabling reliable localization in large-scale street scenes. First, the novel MCL incorporating spatial verification adjusts particle weights by leveraging the rich information embedded in local features, enhancing localization robustness in feature-insufficient scenes by avoiding particles converging to erroneous regions. Second, the pose uncertainty-guided reliable global localization framework monitors the localization status and adaptively switches the localization mode, further improving localization robustness in scenes with insufficient features and incomplete map coverage by exploiting the exploratory capability of particles in MCL. We validated the effectiveness of our proposed method on a large-scale heterogeneous point cloud dataset, which comprises high-precision vehicle-mounted MLS point clouds and helmet-mounted WLS point clouds. The experimental results demonstrate that our method achieves a position accuracy of 1.66m and a yaw accuracy of 3.09 degrees in large-scale street scenes, with an average time cost of approximately 68ms per submap. It exhibits excellent performance in terms of robustness, accuracy, and efficiency. Furthermore, ablation experiments confirm the rationality and effectiveness of the MCL method using spatial verification and the reliable global localization framework guided by pose uncertainty. In the future, we aim to enhance the descriptiveness, discrimination, and generalization of features from place recognition, and to fuse multi-sensor information, such as from IMU, in a tightly coupled manner, to further improve localization performance.
@@ Appendix A Supplementary for quantitative evaluation
Table 6 shows the localization accuracy of Reliable-loc in different localization modes. As shown in the figure, in $PF$ mode, the position accuracy of Reliable-loc ranges from 1.90m to 6.29m, and the yaw accuracy ranges from 3.64 to 7.86 degrees. In $Reg$ mode, the position accuracy of Reliable-loc ranges from 0.62m to 1.67m and the yaw accuracy ranges from 1.95 to 4.39 degrees. The localization accuracy in $Reg$ mode is much higher than that of in $PF$ mode.
@@
width = colspec = Q[87]Q[162]Q[94]Q[106]Q[131]Q[106]Q[125]Q[125], cell21 = r=3, cell51 = r=3, hline1-2,5,8 = -0.08em, Loc Mode & Metrics CS college Info campus Zhongshan park Jiefang road Yanjiang road 1 Yanjiang road 2
PF Rate@2m5deg (%)↑ 36.03 14.86 11.49 25.70 21.58 9.59
Position Error (m)↓ 1.90 ± 2.00 4.48 ± 5.22 6.29 ± 35.23 3.32 ± 4.13 3.10 ± 3.66 5.19 ± 6.21
Yaw Error (deg)↓ 3.64 ± 4.45 5.49 ± 7.66 7.86 ± 10.58 5.33 ± 6.76 4.49 ± 5.78 6.67 ± 8.46
Reg Rate@2m5deg (%)↑ 87.20 71.86 87.56 82.00 61.88 77.68
Position Error (m)↓ 0.62 ± 0.97 1.38 ± 1.71 0.65 ± 1.52 1.67 ± 5.53 1.22 ± 1.79 0.87 ± 1.70
Yaw Error (deg)↓ 1.95 ± 3.45 3.29 ± 5.19 1.94 ± 3.55 2.16 ± 3.87 4.39 ± 6.39 2.80 ± 4.69 Note: The results in the $PF$ localization mode do not count data before MCL converges.
Fig.13 shows the estimated trajectories of all methods on the experimental data. As shown in the figure, Reliable-loc is much more robust than other methods.
Figure 13: Localization trajectories of each method on the experimental data are as follows: the black, blue, green, chocolate, cyan, and red lines represent the trajectories of Ground Truth, PF-loc, PF-SGV-loc, PF-SGV2-loc, Reg-loc, and Reliable-loc, respectively. The green pentagrams indicate the starting points, while the red pentagrams mark the end points. For the convenience of presentation, only the trajectory points with position errors less than 100m are plotted.
@@ Appendix B Supplementary for parameter analysis
Fig.14 shows the localization success rate curves of Reliable-loc in different $\lambda_{thr}$. As shown in the figure, Reliable-loc usually works well when $\lambda_{thr}$ is no more than 1e-3, and its performance exhibits little variation for parameters less than 1e-3.
Figure 14: Localization success rate curves of Reliable-loc with different $\lambda_{thr}$. (a), (b), (c), (g), (h), (i) are the $Rate@Xm$ curves, and (d), (e), (f), (j), (k), (l) are the $Rate@Ydeg$ curves.
Appendix C Supplementary for efficiency analysis
Table 7 shows the average time cost of the key steps of Reliable-loc. As shown in the table, particle clustering accounts for 30.69ms, global feature-based particle weight updating takes 229.54ms, spatial verification-based particle weight adjustment consumes 11.19ms, and localization status monitoring adds 7.12ms.
Table 7: The average time cost of the key steps of Reliable-loc.
@@ Appendix C Supplementary for efficiency analysis
width = colspec = Q[175]Q[150]Q[244]Q[190]Q[177], hlines = 0.08em, Methodological steps & Particle Clustering Global Feature Weight Updating GV Weight Adjustment Loc Status Monitoring
time cost (ms) 30.69 229.54 11.19 7.12
Table 8 shows the average time cost of Reliable-loc in different localization modes. As shown in the table, Reliable-loc requires an average of 1219.69ms and 126.10ms to process one submap before MCL is not converged and converged respectively, and it requires an average of 10.93ms to process one submap in $Reg$ localization mode.
Table 8: The average time cost of Reliable-loc in different localization modes.
@@
width = colspec = Q[210]Q[333]Q[308]Q[85], hlines = 0.08em, Loc mode & PF before convergency PF after convergency Reg
time cost (ms) 1219.69 126.10 10.93

### v2

@@ 6.1 Effectiveness of the spatial verification-based observation model
To further demonstrate the effectiveness of the spatial verification-based observation model, we extracted four 100-frame clips from the experimental data of Jiefang road 1: 500-600, 1000-1100, 1500-1600, and 2000-2100. We then compared the localization performance of PF-loc, PF-SGV-loc, and PF-SGV2-loc on these clips. The particles are initialized with a known position but an unknown yaw in the experiment. The experimental results are presented in Fig.9.
@@ 6.2 Necessity of switching the localization mode
To illustrate the necessity of switching the localization mode, we select representative clips of lengths 750, 500, 500, and 500 frames on the data Info campus, Zhongshan park, Jiefang road 1, and Yanjiang road 1 respectively. We use Reg-loc and Reliable-loc to perform global localization in these clips, the position and yaw angle are known at particle initialization, and the localization trajectories are shown in Fig.10.
@@ 6.3 Distribution of the localization mode
width = colspec = Q[156]Q[90]Q[100]Q[125]Q[112]Q[112]Q[119]Q[119], hlines, Methods & CS college Info campus Zhongshan park Jiefang road 1 Jiefang road 2 Yanjiang road 1 Yanjiang road 2
Reg Mode Ratio (%) 85.61 65.56 90.69 83.5 43.34 36.78 87.24
@@ 6.4 Parameter analysis
As shown in Table 4, Reliable-loc exhibits reduced sensitivity to parameters in Info campus, achieving position and yaw accuracies of approximately ±2.5 m and ±4 degrees, respectively. For CS college, Zhongshan park, Jiefang road 1, Jiefang road 2, and Yanjiang road 2, the optimal parameter is 5e-4. For Yanjiang road 1, the optimal parameter is 1e-4. Reliable-loc demonstrates its ability to achieve reliable localization across various parameter settings, maintaining relatively stable performance on the experimental data. Typically, optimal localization performance is achieved when $\lambda_{thr}$ is set to 1e-4, 5e-4, or 1e-3. In practical applications, a setting of 5e-4 for $\lambda_{thr}$ is recommended.
@@
width = colspec = Q[113]Q[108]Q[83]Q[92]Q[113]Q[102]Q[102]Q[110]Q[110], cell21 = r=2, cell41 = r=2, cell61 = r=2, cell81 = r=2, cell101 = r=2, cell121 = r=2, cell141 = r=2, hline1-2,4,6,8,10,12,14,16 = -, Param (unit:e-4) & Metrics CS college Info campus Zhongshan park Jiefang road 1 Jiefang road 2 Yanjiang road 1 Yanjiang road 2
1 PE (m)↓ 0.87 2.56 1.47 1.92 5.23 2.39 1.51
YE (deg)↓ 2.17 4.19 2.69 2.81 5.23 4.45 3.37
5 PE (m)↓ 0.73 2.41 1.13 1.92 4.93 3.60 1.33
YE (deg)↓ 2.09 4.02 2.44 2.64 4.78 5.62 3.22
10 PE (m)↓ 0.90 2.30 4.14 2.18 4.77 2.89 2.18
YE (deg)↓ 2.53 4.06 3.02 3.24 4.84 4.74 3.16
20 PE (m)↓ 0.91 2.38 1.43 2.21 5.43 3.06 3.21
YE (deg)↓ 2.18 4.20 3.09 3.40 4.91 4.21 4.32
40 PE (m)↓ 0.83 2.22 2.27 2.59 5.83 3.18 4.01
YE (deg)↓ 2.30 3.66 3.73 4.10 5.17 4.23 4.77
80 PE (m)↓ 0.78 2.32 2.32 2.75 6.30 3.18 3.95
YE (deg)↓ 2.49 4.16 3.77 4.06 4.99 4.23 4.55
160 PE (m)↓ 1.36 2.33 2.94 2.76 6.37 3.18 3.95
YE (deg)↓ 2.54 4.07 3.99 4.18 5.33 4.23 4.55
@@ 6.5 Efficiency analysis
As shown in Table 5, Reliable-loc requires an average of 98.04 ms to process one submap. The MCL-based methods exhibit longer total elapsed times, exceeding Reliable-loc by approximately 50 ms. Evidently, Reliable-loc demonstrates superior time efficiency, making it potential for real-time applications.
@@ 6.6 Deficiencies and future work
The above experimental results demonstrate that Reliable-loc can achieve reliable localization in large-scale, complex street scenes by integrating place recognition and MCL, with a position accuracy of up to ±0.73 m and yaw accuracy of up to ±2.09 degrees. Nevertheless, the method’s robustness and accuracy rely on the descriptiveness and discrimination of the global and local features from place recognition. When Reliable-loc starts localization in a scene with extremely insufficient features, the particles are prone to converging to the incorrect region due to the poor discrimination of the global features from place recognition, ultimately resulting in localization failure. A corresponding failure case is shown in Fig.12, and Fig.12 shows the satellite image, helmet point cloud, the curve of the positional error, and the number of grids occupied by the particles over time.
@@
In the future, we will conduct research in four areas to enhance the robustness and accuracy of the localization systems. Firstly, building on existing place recognition efforts, we will introduce contrastive learning to strengthen the descriptiveness, discrimination, and generalization of features extracted by the network, thereby improving the reliability of the localization systems in challenging scenes such as feature insufficiency (Jing and Tian, 2020; Xie et al., 2020; Wu et al., 2023). Secondly, we will explore continuous learning to enable the place recognition model to evolve, overcoming challenges posed by scene switching and sensor changes during long-term localization, and enhancing the system’s resilience (Knights et al., 2022; Yin et al., 2023b). Thirdly, referencing the backend optimization framework in Simultaneous Localization and Mapping, we will fully leverage the information contained in raw data from multiple sensors like IMU and LiDAR to achieve more reliable and accurate localization in a tightly coupled manner (Chiang et al., 2019; Ye et al., 2019; Pan et al., 2021). Finally, repetitive or featureless terrains pose significant challenges to the robustness of localization, and It’s difficult to resolve the problem with only point clouds. We therefore consider integrating non-visual sensors such as IMU and UWB to enhance the robustness of localization in repetitive and featureless terrains.
@@ 7 Conclusion
In this paper, we propose a LiDAR-based sequential global localization method that effectively overcomes challenges posed by feature insufficiency and incomplete map coverage by utilizing spatial and temporal verifiable cues, enabling reliable localization in large-scale street scenes. First, the novel MCL incorporating spatial verification adjusts particle weights by leveraging the rich information embedded in local features, enhancing localization robustness in feature-insufficient scenes by avoiding particles converging to erroneous regions. Second, the pose uncertainty-guided reliable global localization framework monitors the localization status and adaptively switches the localization mode, further improving localization robustness in scenes with insufficient features and incomplete map coverage by exploiting the exploratory capability of particles in MCL. We validated the effectiveness of our proposed method on a large-scale heterogeneous point cloud dataset, which comprises high-precision vehicle-mounted MLS point clouds and helmet-mounted WLS point clouds. The experimental results demonstrate that our method achieves a position accuracy of ±2.91 m and a yaw accuracy of ±3.74 degrees in large-scale street scenes, with an average time cost of approximately 98 ms per submap. It exhibits excellent performance in terms of robustness, accuracy, and efficiency. Furthermore, ablation experiments confirm the rationality and effectiveness of the MCL method using spatial verification and the reliable global localization framework guided by pose uncertainty. In the future, we aim to enhance the descriptiveness, discrimination, and generalization of features from place recognition, and to fuse multi-sensor information, such as from IMU, in a tightly coupled manner, to further improve localization performance.
@@ Appendix A Supplementary for quantitative evaluation
Table 6 shows the localization accuracy of Reliable-loc in different localization modes. As shown in the figure, in $PF$ mode, the position accuracy of Reliable-loc ranges from ±1.90 m to ±7.82 m, and the yaw accuracy ranges from ±3.64 to ±7.86 degrees. In $Reg$ mode, the position accuracy of Reliable-loc ranges from ±0.62 m to ±1.67 m and the yaw accuracy ranges from ±1.95 to ±4.39 degrees. The localization accuracy in $Reg$ mode is much higher than that of in $PF$ mode.
@@
width = colspec = Q[81]Q[115]Q[87]Q[96]Q[117]Q[106]Q[106]Q[113]Q[113], cell21 = r=3, cell51 = r=3, hline1-2,5,8 = -, Loc Mode & Metrics CS college Info campus Zhongshan park Jiefang road 1 Jiefang road 2 Yanjiang road 1 Yanjiang road 2
PF R@2m5d (%)↑ 36.03 14.86 11.49 25.70 0.60 21.58 9.59
PE (m)↓ 1.90 ± 2.00 4.48 ± 5.22 6.29 ± 35.23 3.32 ± 4.13 7.82 ± 8.90 3.10 ± 3.66 5.19 ± 6.21
YE (deg)↓ 3.64 ± 4.45 5.49 ± 7.66 7.86 ± 10.58 5.33 ± 6.76 5.80 ± 8.20 4.49 ± 5.78 6.67 ± 8.46
Reg R@2m5d (%)↑ 87.20 71.86 87.56 82.00 70.34 61.88 77.68
PE (m)↓ 0.62 ± 0.97 1.38 ± 1.71 0.65 ± 1.52 1.67 ± 5.53 1.18 ± 1.84 1.22 ± 1.79 0.87 ± 1.70
YE (deg)↓ 1.95 ± 3.45 3.29 ± 5.19 1.94 ± 3.55 2.16 ± 3.87 3.45 ± 5.19 4.39 ± 6.39 2.80 ± 4.69 Note: The results in the $PF$ localization mode do not count data before MCL converges.
Fig.13 shows the PVA (position, velocity, and angle) errors of the proposed method on the experimental data. As shown in the figure, Reliable-loc performs well in data of easy and medium difficulty, such as CS college and Zhongshan park. In data of hard difficulty, such as Info campus and Jiefang road 2, the lack of features in the scene, data holes in the map, and changes in LiDAR sensor types have negative impacts on the localization accuracy of the proposed method.
Figure 13: PVA errors of the proposed method. The black color represents the velocity of the WLS at different time points, while the green, red, and blue colors represent the localization errors of the proposed method at different time points, respectively.
@@ Appendix B Supplementary for parameter analysis
Besides, four 500-frame clips, 0-500 and 500-100 from CS college, 500-1000, 1500-2000 from Jiefang road 1, are used to analyze the impact of the number of particles on localization during initialization and after the convergence of particle filtering. Table 7 shows the localization accuracy and time cost of Reliable-loc with different numbers of initial particles. As shown in Table 7, Reliable-loc performs well in CS college with no less than 500 initial particles and it performs well in Jiefang road 1 with no less than 1000 initial particles. Its time cost rapidly increases with the number of initial particles. Table 8 shows the localization accuracy of Reliable-loc with different numbers of particles after the convergence of particle filtering. As shown in Table 8, the number of particles after the convergency of particle filtering has a minor impact on localization performance, and this paper selects a relatively large number of particles to ensure localization robustness.
Table 7: Localization accuracy and time cost of Reliable-loc with different numbers of initial particles.
@@
width = colspec = Q[106]Q[154]Q[127]Q[127]Q[106]Q[106]Q[106]Q[106], cell21 = r=3, cell51 = r=3, cell81 = r=3, cell111 = r=3, hline1-2,5,8,11,14 = -, Data clips & Num of
initial particles
100 500 1000 2500 5000 10000
1 PE (m)↓ 44.41 ± 48.16 8.15 ± 115.84 0.84 ± 6.34 0.61 ± 1.55 0.79 ± 2.28 0.57 ± 0.95
YE (deg)↓ 22.98 ± 28.01 2.93 ± 13.18 2.51 ± 7.48 2.38 ± 8.72 2.45 ± 4.63 2.07 ± 3.68
time cost (ms)↓ 132.57 117.4 123.03 162.28 369.44 702.1
2 PE (m)↓ 46.07 ± 74.63 19.38 ± 36.84 0.52 ± 1.00 0.65 ± 1.35 0.64 ± 1.19 0.48 ± 0.78
YE (deg)↓ 35.17 ± 56.20 16.85 ± 32.95 2.04 ± 3.53 2.60 ± 6.48 2.36 ± 4.48 1.85 ± 3.18
time cost (ms)↓ 112.44 158.93 125.39 172.32 271.7 659.73
3 PE (m)↓ 1.81 ± 3.37 0.85 ± 1.58 1.02 ± 1.54 0.74 ± 1.19 1.33 ± 2.00 1.10 ± 1.66
YE (deg)↓ 3.83 ± 7.82 2.33 ± 4.04 2.58 ± 4.48 2.37 ± 4.11 4.33 ± 6.93 3.35 ± 5.48
time cost (ms)↓ 112.59 118 133.38 171.72 440.26 474.28
4 PE (m)↓ 46.73 ± 48.44 1.13 ± 1.66 1.50 ± 3.21 1.35 ± 2.10 1.85 ± 3.46 0.99 ± 2.38
YE (deg)↓ 23.14 ± 32.31 2.60 ± 4.11 3.38 ± 8.68 2.58 ± 4.06 4.05 ± 8.80 2.71 ± 5.70
time cost (ms)↓ 129.41 152.94 149.59 194.51 220.04 459.72
Table 8: Localization accuracy of Reliable-loc with different numbers of particles after the convergence of particle filtering.
@@
width = colspec = Q[98]Q[171]Q[113]Q[113]Q[113]Q[113]Q[104]Q[104], cell21 = r=2, cell41 = r=2, cell61 = r=2, cell81 = r=2, hline1-2,4,6,8,10 = -, Data clips & Num of particles
after convergence
25 50 100 200 400 800
1 PE (m)↓ 0.86 ± 2.38 0.76 ± 2.32 0.77 ± 2.35 0.83 ± 2.41 0.87 ± 2.41 0.90 ± 2.58
YE (deg)↓ 2.52 ± 4.84 2.18 ± 4.39 2.27 ± 4.54 2.14 ± 4.16 2.53 ± 5.07 2.28 ± 4.32
2 PE (m)↓ 0.50 ± 0.84 0.59 ± 1.06 0.54 ± 0.95 0.58 ± 1.10 0.66 ± 1.22 0.61 ± 1.14
YE (deg)↓ 1.78 ± 3.24 2.08 ± 3.89 1.87 ± 3.53 2.12 ± 3.99 2.36 ± 4.71 2.01 ± 3.70
3 PE (m)↓ 1.59 ± 2.47 0.99 ± 1.44 0.93 ± 1.39 1.27 ± 2.21 1.18 ± 1.70 1.18 ± 1.76
YE (deg)↓ 3.62 ± 6.35 3.25 ± 5.50 3.13 ± 4.95 3.95 ± 6.20 3.49 ± 5.56 3.63 ± 5.57
4 PE (m)↓ 1.54 ± 3.65 1.28 ± 3.50 2.27 ± 4.16 1.85 ± 3.94 1.35 ± 2.94 2.14 ± 3.74
YE (deg)↓ 5.47 ± 16.68 5.03 ± 16.03 5.46 ± 15.91 5.33 ± 15.86 3.79 ± 8.47 3.93 ± 7.13
Appendix C Localization robustness to partial observations, dynamic objects, and content changes
Although partial observations, dynamic objects, and content changes are not the primary focus of this paper, we still select four 200-frame clips from Jiefang road 2 and conduct additional experiments refer to section 6.1 to demonstrate the robustness of our method against these challenges. As shown in Fig.14, the localization performance of PF-loc, PF-SGV-loc, PF-SGV2-loc, and Reliable-loc is improved sequentially on all four clips. In clip 1, lots of pedestrians and vehicles pose challenges to localization. In clip 2, occlusions from the long closure limits the MLS and WLS systems to capturing only very limited scene content. In clip 3 and 4, due to a time difference of over 4 years between the MLS and WLS point cloud acquisitions, trees in some scenes have grown significantly, and construction closures have been newly added in other scenes. Owing to the proposed spatial verification-based MCL and localization status monitoring mechanism, Reliable-loc performs well in all four clips.
Figure 14: Localization trajectories in the four clips when the initial position is known and the yaw is unknown. The black, blue, green, orange, cyan, and red lines are the trajectories of the Ground Truth, PF-loc, PF-SGV-loc, PF-SGV2-loc, Reg-loc, and Reliable-loc, respectively, with the green pentagram as the starting point and the red pentagram as the end point.